how do you graph y=1/3x-2
Answer by checkley77(7072)

y=1/3x-2
y=x/3-2 This line has a slope=1/3 & a y intercept=-2.
(graph 300x200 pixels, x from -6 to 7, y from -10 to 10, x/3 -2).
(6,0), (0,-2) are the x & y intercepts.